# The Anderson transition – a view from Krylov space

### v2

@@ I Introduction
The edge mode nature of LIOMs studied below has been identified in recent work on nearly integrable spin chains [Yates2020_LifetimeofAlmostStrong, Yates2021StrongandAlmoststrongModes, Yeh2023SlowlyDecayingZeroMode]. By way of surveying other recent efforts to exploit Krylov methods, we note there have been many studies of quantum chaos and integrability in Krylov space [Parker2019_universaloperatorgrowth, Barbón2019OnTheEvolutionofOperatorComplexity, Rabinovici2021_OperatorComplexityaJourney, Rabinovici2022Krylovlocalizationandsuppression, Hornedal2022_UltimateSpeedLimitstoGrowthofOpComplexity, Caputa2022_GeometryofKrylovComplexity, Balasubramanian2022_QuantumChaosandtheComplexityofSpreadofStates, Bhattacharjee2022_KComplexity_SaddleDominatedScrambling, Erdmenger2023_UniversalChaoticDynamicsfromKrylov, Rabinovici2022_KrylovComplexityfromIntegrabilitytoChaos, Hashimoto2023_KrylovComplexityandChaos, Suchsland2025_KrylovComplexityandTrotterTransitions, Camargo2024_SpectralandKrylovComplexityBilliards, Cao_2021StatisticalmethodforOG, menzler2024krylovlocalizationprobeergodicity, baggioli2024_krylovcomplexityorderparameter, Yates2020DynamicsofAlmostStrongEdgeModes, Yates2021StrongandAlmoststrongModes, Yeh2023SlowlyDecayingZeroMode, Yeh2024UniversalModelofFloquet, loizeau2025openingkrylovspaceaccess, Trigueros2022_KrylovComplexityofMBL, Balasubramanian2025_ChaosIntegrabilityandLateTimes, Yates2020_LifetimeofAlmostStrong, Scialchi2025_ExploringQuantumErgodicity_KrylovApproach, Uskov2024_QuantumDynamicsinOneandTwoDims, Teretenkov_2025_PseudomodeExpansion, Loizeau_2025_PauliStrings, Yeh2025_MomentMethodandContinuedFractionExpansion] which come with various proposals for detecting ergodicity breaking transitions. For one, chaos has been associated with a maximum growth rate of the Lanczos coefficients [Parker2019_universaloperatorgrowth], though this is not always the case [Bhattacharjee2022_KComplexity_SaddleDominatedScrambling]. The interplay between integrability, integrability breaking, and dimerization has been pointed out in Refs. [Rabinovici2022Krylovlocalizationandsuppression, Hashimoto2023_KrylovComplexityandChaos, nandy2024quantumdynamicskrylovspace, menzler2024krylovlocalizationprobeergodicity]. Finally, chaos is associated with delocalization in Krylov space [Dymarsky2020_QuantumChaosAsDelocalization], which can be quantified by the average position in Krylov space known as the “Krylov complexity”, the long time average of which is expected to saturate a maximum bound for chaotic systems [Parker2019_universaloperatorgrowth, Barbón2019OnTheEvolutionofOperatorComplexity, Rabinovici2021_OperatorComplexityaJourney, Rabinovici2022Krylovlocalizationandsuppression, Hornedal2022_UltimateSpeedLimitstoGrowthofOpComplexity, Caputa2022_GeometryofKrylovComplexity, Balasubramanian2022_QuantumChaosandtheComplexityofSpreadofStates, Bhattacharjee2022_KComplexity_SaddleDominatedScrambling, Erdmenger2023_UniversalChaoticDynamicsfromKrylov, Rabinovici2022_KrylovComplexityfromIntegrabilitytoChaos, Hashimoto2023_KrylovComplexityandChaos, Camargo2024_SpectralandKrylovComplexityBilliards].
@@ III.2 Critical Disorder
The Krylov space properties of the critical point can be constrained by the same analysis we’ve conducted on the delocalized side. The Anderson critical point is known to exhibit multi-fractality [Janssen1994_Multifractal_localization, Evers_2008_AndersonTransitions] and was recently studied in detail in Ref. [Hopjan2023_Scale_InvariantSurvivalProb]. This multi-fractality manifests itself as powerlaw scaling of the survival probability $C(t)\sim t^{-\beta}$, with $\beta\approx 0.42$. At the critical point, the Thouless time approaches the Heisenberg time (which is $O(1/V)$, implying
@@ IV Conclusions and Outlook
In summary, we have described the form of LIOMs in the Anderson model in Krylov space as generated by projectors on single sites. There are three types of behaviors – stretched exponential decay inside the localized phase (with geometric exponent $\gamma=1/2d$ set by real space dimension $d$), delocalization on real-space volume $V=\sqrt{D_{K}}$ states in the metallic/extended phase, and powerlaw decay at the critical point. Importantly, these results may be interpreted using effective disorder-free chains obtained from disorder averaged Krylov coefficients, which allows considerable further insight into the Anderson transition: e.g. the significance of Krylov dimerization as a mechanism of localizing operators despite having a gapless Liouvillian.
Our work may be extended to other aspects of the Anderson transition: for example, to perhaps enable the tracking of the mobility edge as a function of typical energy. It also sets up a framework for extending these results to interacting models where the coefficients grow with some power $b_{n}\sim n^{\delta}$ (with $\delta\leq 1$). If the dimerization decays like a powerlaw as in our model $\sim n^{-(1-\gamma)}$ then – since only the ratio of Lanczos coefficients matters for the zero-mode – the growth simply re-normalizes the power to be $\sim n^{-(1-\gamma-\delta)}$. Consequently, the faster the coefficients grow, i.e. the faster the operator growth, the slower the dimerization must decay in order for conserved quantities to exist. For spin systems debated to show MBL, such as the disordered Heisenberg model, the coefficients have been found to grow as $\sim n/\log{n}$ on average, with some dimerization [Trigueros2022_KrylovComplexityofMBL, Weisse2025_OperatorGrowthMBL]. The $\log$ correction would only be relevant at the critical power implying that the dimerization has to be constant for conserved quantities to exist (i.e. $\gamma=1$, its maximum allowed value). Additionally, it would be interesting to understand how the dimerization behaves in interacting integrable models such as the XXZ-chain, where there is typically square root growth ($\delta=0.5$) and thus only $1\geq\gamma\geq 0.5$ is allowed to host the conserved quantities we know to exist.

### v3

@@ I Introduction
The edge mode nature of LIOMs studied below has been identified in recent work on nearly integrable spin chains [Yates2020_LifetimeofAlmostStrong, Yates2021StrongandAlmoststrongModes, Yeh2023SlowlyDecayingZeroMode]. By way of surveying other recent efforts to exploit Krylov methods, we note there have been many studies of quantum chaos and integrability in Krylov space [Parker2019_universaloperatorgrowth, Barbón2019OnTheEvolutionofOperatorComplexity, Rabinovici2021_OperatorComplexityaJourney, Rabinovici2022Krylovlocalizationandsuppression, Hornedal2022_UltimateSpeedLimitstoGrowthofOpComplexity, Caputa2022_GeometryofKrylovComplexity, Balasubramanian2022_QuantumChaosandtheComplexityofSpreadofStates, Bhattacharjee2022_KComplexity_SaddleDominatedScrambling, Erdmenger2023_UniversalChaoticDynamicsfromKrylov, Rabinovici2022_KrylovComplexityfromIntegrabilitytoChaos, Hashimoto2023_KrylovComplexityandChaos, Suchsland2025_KrylovComplexityandTrotterTransitions, Camargo2024_SpectralandKrylovComplexityBilliards, Cao_2021StatisticalmethodforOG, menzler2024krylovlocalizationprobeergodicity, baggioli2024_krylovcomplexityorderparameter, Yates2020DynamicsofAlmostStrongEdgeModes, Yates2021StrongandAlmoststrongModes, Yeh2023SlowlyDecayingZeroMode, Yeh2024UniversalModelofFloquet, loizeau2025openingkrylovspaceaccess, Trigueros2022_KrylovComplexityofMBL, Balasubramanian2025_ChaosIntegrabilityandLateTimes, Yates2020_LifetimeofAlmostStrong, Scialchi2025_ExploringQuantumErgodicity_KrylovApproach, Uskov2024_QuantumDynamicsinOneandTwoDims, Teretenkov_2025_PseudomodeExpansion, Loizeau_2025_PauliStrings, Yeh2025_MomentMethodandContinuedFractionExpansion, Bartsch2024, angelinos2025temperaturedependencekrylovspace] which come with various proposals for detecting ergodicity breaking transitions. For one, chaos has been associated with a maximum growth rate of the Lanczos coefficients [Parker2019_universaloperatorgrowth], though this is not always the case [Bhattacharjee2022_KComplexity_SaddleDominatedScrambling]. The interplay between integrability, integrability breaking, and dimerization has been pointed out in Refs. [Rabinovici2022Krylovlocalizationandsuppression, Hashimoto2023_KrylovComplexityandChaos, nandy2024quantumdynamicskrylovspace, menzler2024krylovlocalizationprobeergodicity]. Finally, chaos is associated with delocalization in Krylov space [Dymarsky2020_QuantumChaosAsDelocalization], which can be quantified by the average position in Krylov space known as the “Krylov complexity”, the long time average of which is expected to saturate a maximum bound for chaotic systems [Parker2019_universaloperatorgrowth, Barbón2019OnTheEvolutionofOperatorComplexity, Rabinovici2021_OperatorComplexityaJourney, Rabinovici2022Krylovlocalizationandsuppression, Hornedal2022_UltimateSpeedLimitstoGrowthofOpComplexity, Caputa2022_GeometryofKrylovComplexity, Balasubramanian2022_QuantumChaosandtheComplexityofSpreadofStates, Bhattacharjee2022_KComplexity_SaddleDominatedScrambling, Erdmenger2023_UniversalChaoticDynamicsfromKrylov, Rabinovici2022_KrylovComplexityfromIntegrabilitytoChaos, Hashimoto2023_KrylovComplexityandChaos, Camargo2024_SpectralandKrylovComplexityBilliards].
@@ III.2 Critical Disorder
The Krylov space properties of the critical point can be constrained by the same analysis we’ve conducted on the delocalized side. The Anderson critical point is known to exhibit multi-fractality [Janssen1994_Multifractal_localization, Evers_2008_AndersonTransitions] and was recently studied in detail in Ref. [Hopjan2023_Scale_InvariantSurvivalProb]. This multi-fractality manifests itself as powerlaw scaling of the survival probability $C(t)\sim t^{-\beta}$, with $\beta\approx 0.42$. At the critical point, the Thouless time approaches the Heisenberg time (and is $O(1/V)$) implying
@@ IV Conclusions and Outlook
In summary, we have described the form of LIOMs in the Anderson model in Krylov space as generated by projectors on single sites. There are three types of behaviors – stretched exponential decay inside the localized phase (with geometric exponent $\gamma=1/2d$ set by real space dimension $d$), delocalization on real-space volume $V=\sqrt{D_{K}}$ states in the metallic/extended phase, and powerlaw decay at the critical point. Importantly, these results may be interpreted using effective disorder-free chains obtained from disorder averaged Krylov coefficients, which allows considerable further insight into the Anderson transition: e.g. the significance of Krylov dimerization as a mechanism of localizing operators despite having a gapless Liouvillian. The critical point occurs when the average dimerization decreases faster than a powerlaw with power 1, which is the point at which the zero-modes cease to be normalizable in Krylov space ($\gamma=0$ in expression (10)).
Our work may be extended to other aspects of the Anderson transition: for example, to perhaps enable the tracking of the mobility edge as a function of typical energy. It also sets up a framework for extending these results to interacting models where the coefficients grow with some power $b_{n}\sim n^{\delta}$ (with $\delta\leq 1$). If the dimerization decays like a powerlaw – since only the ratio of Lanczos coefficients matters for the zero-mode – the growth simply re-normalizes that power. Consequently, the faster the coefficients grow, i.e. the faster the operator growth, the slower the dimerization must decay in order for conserved quantities to exist. For spin systems debated to show MBL, such as the disordered Heisenberg model, the coefficients have been found to grow as $\sim n/\log{n}$ on average, with some dimerization [Trigueros2022_KrylovComplexityofMBL, Weisse2025_OperatorGrowthMBL]. The $\log$ correction would only be relevant at the critical power implying that the dimerization has to be constant for conserved quantities to exist. Interestingly, this still leaves the possibility for the dimerization to decay in interacting integrable models such as the Heisenberg model, where there is typically square root growth [Parker2019_universaloperatorgrowth].